# Planet formation regulated by galactic-scale interstellar turbulence

### v2

@@ 3.5 Disc turbulence and angular momentum transport:
To estimate the degree of ISM-induced turbulence, we must compute the balance between the rate at which energy is deposited compared to the decay of that turbulent energy. The rate at which energy decays is (Klessen & Hennebelle, 2010):
$\dot{E}_{\mathrm{decay}}=-\frac{1}{2}\frac{M_{\mathrm{disc}}v_{\mathrm{t,disc}}^{3}}{H_{\mathrm{disc}}},$ (12)
@@
where $H_{\mathrm{disc}}\sim 0.1R_{\mathrm{disc}}$ is the disc scale height. The in-flowing turbulent energy from the ISM is:
@@
Therefore, the BHL accretion rate $\dot{M}_{\mathrm{BHL}}$ needs to remain sufficient to drive observed turbulence throughout the disc lifetime. However, this requirement is probably met for most discs in the context of our model. With the exception of DM Tau, which has turbulent velocity $v_{\rm{t,disc}}\sim 0.25{-}0.33\,c_{\mathrm{s}}$, and IM Lup which has $v_{\rm{t,disc}}\sim 0.4{-}0.6\,c_{\mathrm{s}}$ (Paneque-Carreño et al., 2024) attempts to measure turbulence in protoplanetary discs have resulted in upper limits. For example, for the disc around HD 163296 this upper limit is $v_{\rm{t,disc}}\lesssim 0.05\,c_{\mathrm{s}}$ (Flaherty et al., 2015, 2017) and for TW Hydra it is $v_{\rm{t,disc}}\lesssim 0.08\,c_{\mathrm{s}}$ (Flaherty et al., 2018; Teague et al., 2018). For $M_{\mathrm{disc}}/m_{*}=0.01$, $H_{\mathrm{disc}}/R_{\mathrm{disc}}=0.1$ and $\epsilon_{\rm{t}}=1$, then even substituting a modest $\dot{M}_{\rm BHL}\sim 10^{-11}\,M_{\odot}$ yr-1 yields $v_{\mathrm{t,disc}}\sim 0.03c_{\mathrm{s}}$, similar to the empirical upper limit on turbulence in some discs. This is approximately the median BHL accretion rate for a solar mass star in our model at an age of $\sim 3$ Myr (see the bottom panel of Figure 2). Typically values of $M_{\mathrm{disc}}$ are even smaller than $10^{-2}m_{*}$, so even greater turbulence may be imparted by the ISM. Broadly, for discs older than $\sim 3$ Myr (as is the case for all those with upper limits on the turbulence) we expect relatively low levels of turbulence, while for younger discs (like DM Tau (Zhang et al., 2020) and IM Lup (Mawet et al., 2012)) turbulence driven by the ISM can also be larger.
@@
although in practice $|\dot{E}_{\mathrm{acc}}|\ll|\dot{E}_{\mathrm{decay}}|$. Then we can solve the initial value problem, including the mass evolution of the disc as before. In principle, we could then compute the disc evolution fully consistently assuming only the ISM as the driver of internal physics, since we then have $\alpha_{\mathrm{SS}}$ or equivalently $\tau_{\mathrm{acc}}$. However, in practice, this requires introducing a free parameter $\epsilon_{\mathrm{t}}$, which may not be a constant. Strictly, it also means self-consistently evolving the disc outer radius (or disc scale radius), which should viscously expand, such that we can compute the correct $v_{\mathrm{in}}$. This would also be balanced by the contraction due to deposition of low angular momentum material (Wijnen et al., 2017), resulting in additional (non-turbulent) stellar accretion. For these reasons, self-consistently solving the disc evolution would considerably complicate our model. We therefore leave this to future work. Nonetheless, we can adopt sensible parameters: $\epsilon_{\mathrm{t}}=1$, $R_{\mathrm{disc}}=250(m_{*}/1\,M_{\odot})$ au and $H_{\mathrm{disc}}=0.1R_{\mathrm{disc}}$ with our fixed $\tau_{\mathrm{acc}}$ to obtain an estimate for the maximal turbulent disc velocity $v_{\mathrm{t,disc}}$ over time. The outcome of this exercise is shown in Figure 11, and discussed in the main article. We find that sufficient turbulent energy is supplied to discs from the ISM to sustain the observed turbulence.
@@
We can also consider the specific case of DM Tau, which is a star of mass $\sim 0.4{-}0.5\,M_{\odot}$ hosting a disc with a dust cavity and accreting at a rate $\sim 10^{-8}\,M_{\odot}$ yr-1 (Manara et al., 2014; Francis & van der Marel, 2020). It has a turbulent velocity $v_{\rm{t,disc}}\sim 0.25{-}0.33\,c_{\mathrm{s}}$, or viscous $\alpha_{\mathrm{SS}}\sim 0.1$ (Flaherty et al., 2020). For this disc, a recent and relatively rapid period of accretion would be needed to drive the observed turbulence. Estimating $M_{\mathrm{disc}}/m_{*}\approx 0.03$ (Andrews et al., 2013) this would imply recent accretion from the ISM at a rate $\dot{M}_{\mathrm{BHL}}\gtrsim 2\times 10^{-8}\,M_{\odot}$ yr-1 that ended in the last $\sim 10^{4}$ years. This is similar to the current stellar accretion rate (Manara et al., 2014), suggesting that the majority of the mass of the present day disc could have originated from this putative in-fall event. This scenario is plausible statistically within the context of our model. However, apart from the argument above, there is currently no particular evidence that DM Tau has undergone recent in-fall. This underlines the importance of a systematic search for observational signatures of late stage in-fall events.
3.6 Observability of in-fall events:
@@
where $\epsilon_{\mathrm{geo}}\leq 1$ is a geometric factor accounting for the non-uniform distribution of material and $\mu$ is the mean molecular mass.
@@
The typical column density implied by equation 21 is relatively low, such that we would expect that in-fall is often challenging to trace observationally. We may consider the example of CO isotopologues, which are often the brightest tracers used to probe the gas content of protoplanetary discs with ALMA. These isotopologues have a self-shielding column density of $\sim 10^{15}$ cm-2 (van Dishoeck & Black, 1988; Visser et al., 2009), and typical abundance relative to molecular hydrogen $\sim 10^{-5}$ (Trapman et al., 2023). As an order of magnitude estimate, we may therefore expect column densities $\lesssim 10^{20}$ cm-2 to have the CO photodissociated. From equation 21, we plot the $N_{\mathrm{ext}}$ distribution over time resulting from the accretion rates in our model in Figure 12. We see immediately in-falling CO would be photodissociated for the majority of discs in the age range $\sim 1{-}3$ Myr. Interestingly, in this age range, the fraction of discs above the threshold $N_{\mathrm{ext}}=10^{20}$ cm-2, approximately the CO self-shielding column, drops from approximately $30$ percent to close to 00 percent. Given that this is the range of ages in Taurus, this would be commensurate with the inferred $\sim 16$ percent of discs with evidence of interaction with the ambient medium from scattered light observations (Garufi et al., 2024). However, the physical column density down to which these observations are sensitive is not clear. While these estimates are approximate, they highlight the importance of deep surveys of the environments surrounding large numbers of discs. Such data could be applied to make statistical inferences about the prevalence of late stage in-fall, testing the theoretical predictions we have made in this work.
4 Discussion
4.1 Caveats for the ISM accretion rate:
4.1.1 BHL accretion mediated by a disc
A further pertinent question is what we expect the appearance of the resultant disc to be. As previously discussed, when accreting material with with a high vorticity, the resultant disc outer radius can be of order the BHL radius (Krumholz et al., 2006; Kuffmeier et al., 2023). What is more uncertain is what should be the structure of a disc if the angular momentum of the accreted material is not aligned with the existing material. Depending on the relative angular momentum vector, this could result in ‘puffed up’ discs, or misalignment between the inner and outer regions (Kuffmeier et al., 2021). By contrast, at least the mm-dust component of discs is frequently observed to be extremely flat (Duchêne et al., 2003; Pinte et al., 2007, 2016; Villenave et al., 2020, 2022). However, mm-dust is also commonly more centrally concentrated than the gaseous disc by a factor $\sim 2.5$ (Ansdell et al., 2018; Facchini et al., 2019; Andrews, 2020). Since dust radial drift, growth and vertical settling should occur simultaneously, it is not clear how accretion of turbulent gas in the outer disc should be reflected in the dust distribution. This represents an important topic for future study.
4.1.2 Cloudlet capture
Another possible avenue for the capture of interstellar gas is the capture of small cloudlets via tidal energy dissipation. During the close passage between a star and a cloud, the orbital energy may be dissipated by non-radial oscillations within the cloud, analogous to the formation of tight binaries in globular clusters (Robe, 1968; Fabian et al., 1975), in a process that remains similar across a range of internal density profiles (Lee & Ostriker, 1986). Capture requires a periastron $a_{\mathrm{p}}$ during an encounter between a cloud with radius $R_{\mathrm{cloud}}$ and mass $M_{\mathrm{cloud}}$ and star with mass $m_{*}$ such that:
${a_{\mathrm{p}}}\lesssim{a_{\mathrm{capt}}}\approx{R_{\mathrm{cloud}}}\left[\frac{Gm_{*}}{R_{\mathrm{cloud}}v_{\infty}^{2}}q_{\mathrm{cloud}}^{-1}(1+q_{\mathrm{cloud}})\right]^{1/6}.$ (22)
$\sigma_{\mathrm{capt}}\approx\frac{2\pi Gm_{*}(1+q_{\mathrm{cloud}})\left(a_{\mathrm{capt}}-R_{\mathrm{cloud}}\right)}{v_{\infty}^{2}}.$ (23)
The cloud must be compact enough to be captured fully by the star. Therefore the total rate of capture for a cloud of radius $R_{\mathrm{cloud}}$ is:
$\Gamma_{\mathrm{capt}}(R_{\mathrm{cloud}}|\rho_{\mathrm{cloud}},\sigma_{v},)=\int_{0}^{v_{\mathrm{BHL}}}n_{\mathrm{clouds}}(R_{\mathrm{cloud}})\sigma_{\mathrm{capt}}\tilde{v}g_{3}(\tilde{v}|\sigma_{v})\,\mathrm{d}\tilde{v},$ (24)
where
$g_{N_{\mathrm{D}}}(\Delta v|\sigma_{v})=\frac{2\Delta v^{N_{\mathrm{D}}-1}}{\Gamma(N_{\mathrm{D}}/2)}\left(\frac{1}{4\sigma_{v}^{2}}\right)^{{N_{\mathrm{D}}}/2}\exp\left(\frac{-\Delta v^{2}}{4\sigma_{v}^{2}}\right)$ (25)
is the Maxwell-Boltzmann distributions in $N_{\mathrm{D}}$ dimensions. We have cast equation 24 in terms of the cloud density $\rho_{\mathrm{cloud}}=3M_{\mathrm{cloud}}/4\pi R_{\mathrm{cloud}}^{3}$. If large scale turbulence dominates density fluctuations, we can estimate that a typical cloud density will be $\rho_{\mathrm{cloud}}\sim\rho_{\mathrm{gas}}$. We have introduced $n_{\mathrm{clouds}}=1/(2R_{\mathrm{cloud}})^{3}$, where $2R_{\mathrm{cloud}}$ is the wavelength of the turbulent oscillations in the ISM that produce a cloud of radius $R_{\mathrm{cloud}}$. Finally, we have defined the BHL velocity $v_{\mathrm{BHL}}$, which is the velocity that yields a particular $R_{\mathrm{BHL}}=R_{\mathrm{cloud}}$ when substituted into equation 4. The overall accretion rate is then $\dot{M}_{\mathrm{capt}}=\Gamma_{\mathrm{capt}}M_{\mathrm{cloud}}$.
[…]
4.1.3 Face-on accretion

### v3

@@ 4.1.1 Angular momentum transport to the inner disc
A second way in which BHL accretion may contribute to stellar accretion throughout the disc lifetime is via injecting turbulent energy directly into the disc. We can estimate how much turbulent energy can be supplied to the outer disc by the ISM during BHL accretion by computing the balance between the rate at which energy is deposited compared to the decay of that turbulent energy. The rate at which turbulent energy decays is (Klessen & Hennebelle, 2010):
$\dot{E}_{\mathrm{decay}}=-\frac{3^{3/2}}{2}\frac{M_{\mathrm{disc}}v_{\mathrm{t,disc}}^{3}}{H_{\mathrm{disc}}},$ (10)
@@
where $H_{\mathrm{disc}}\sim 0.1R_{\mathrm{disc}}$ is the disc scale height and the factor $3^{3/2}$ is included because $v_{\mathrm{t,disc}}$ is defined to be the turbulent velocity in a single dimension. The in-flowing turbulent energy from the ISM is:
@@
Therefore, the BHL accretion rate $\dot{M}_{\mathrm{BHL}}$ needs to remain continuously sufficient to drive any observed turbulence throughout the disc lifetime. However, this requirement is probably met for most discs in the context of our model.
@@
although in practice $|\dot{E}_{\mathrm{acc}}|\ll|\dot{E}_{\mathrm{decay}}|$. Then we can solve the initial value problem, including the mass evolution of the disc as before. In principle, we could then compute the disc evolution fully consistently assuming only the ISM as the driver of internal physics, since we then have $\alpha_{\mathrm{SS}}$ or equivalently $\tau_{\mathrm{acc}}$. However, in practice, this requires introducing a free parameter $\epsilon_{\mathrm{t}}$, which may not be a constant. Strictly, it also means self-consistently evolving the disc surface density (or outer radius). For these reasons, self-consistently solving the disc evolution would considerably complicate our model. We therefore leave this to future work. Nonetheless, we can adopt sensible parameters: $\epsilon_{\mathrm{t}}=1$, $R_{\mathrm{disc}}=250(m_{*}/1\,M_{\odot})$ au and $H_{\mathrm{disc}}=0.1R_{\mathrm{disc}}$ with our fixed $\tau_{\mathrm{acc}}$ to obtain an estimate for the maximal turbulent disc velocity $v_{\mathrm{t,disc}}$ over time.
The outcome of this exercise is shown in Figure 8, where we find a median $\alpha_{\mathrm{SS}}\sim 3\times 10^{-3}$ for $\sim 3$ Myr old discs. We can compare this to the $\tau_{\mathrm{acc}}$ distribution (Figure 5), assuming that:
$\tau_{\mathrm{acc}}\approx 3.8\left(\frac{\alpha_{\mathrm{SS}}}{10^{-3}}\right)^{-1}\,\left(\frac{H_{\mathrm{s}}/R_{\mathrm{s}}}{0.1}\right)^{-2}\,\left(\frac{m_{*}}{1\,M_{\odot}}\right)^{-1/2}\,\left(\frac{R_{\mathrm{s}}}{50\,\mathrm{au}}\right)^{3/2}\,\mathrm{Myr},$ (18)
where we define $H_{\mathrm{s}}$ as the scale height at the viscous scale radius $R_{\mathrm{s}}$, which is some fraction of the outer disc radius. Given the median $\tau_{\mathrm{acc}}$ in our model at $3$ Myr is approximately $1$ Myr, the median $\alpha_{\mathrm{SS}}$ is approximately sufficient to sustain accretion self-consistently in our model by direct kinetic energy injection alone (depending on disc geometry and injection efficiency). In Figure 8 we also compare to observational constraints for turbulence in the outer regions of some discs. With the exception of DM Tau, which has turbulent velocity $v_{\rm{t,disc}}\sim 0.25{-}0.33\,c_{\mathrm{s}}$, and IM Lup which has $v_{\rm{t,disc}}\sim 0.4{-}0.6\,c_{\mathrm{s}}$ (Paneque-Carreño et al., 2024) attempts to measure turbulence in protoplanetary discs have resulted in upper limits. For example, for the disc around HD 163296 this upper limit is $v_{\rm{t,disc}}\lesssim 0.05\,c_{\mathrm{s}}$ (Flaherty et al., 2015, 2017) and for TW Hydra it is $v_{\rm{t,disc}}\lesssim 0.08\,c_{\mathrm{s}}$ (Flaherty et al., 2018; Teague et al., 2018). For $M_{\mathrm{disc}}/m_{*}=0.01$, $H_{\mathrm{disc}}/R_{\mathrm{disc}}=0.1$ and $\epsilon_{\rm{t}}=1$, then even substituting a modest $\dot{M}_{\rm BHL}\sim 10^{-11}\,M_{\odot}$ yr-1 yields $v_{\mathrm{t,disc}}\sim 0.03c_{\mathrm{s}}$, similar to the empirical upper limit on turbulence in some discs. This is approximately the median BHL accretion rate for a solar mass star in our model at an age of $\sim 3$ Myr (see the bottom panel of Figure 2). Typically values of $M_{\mathrm{disc}}$ are even smaller than $10^{-2}m_{*}$, so even greater turbulence may be imparted by the ISM. Broadly, for discs older than $\sim 3$ Myr (as is the case for all those with upper limits on the turbulence) we expect relatively low levels of turbulence, while for younger discs, like DM Tau (Zhang et al., 2020) and IM Lup (Mawet et al., 2012), turbulence driven by the ISM can also be larger. We conclude that sufficient turbulent energy is supplied to discs from the ISM to sustain the observational constraints on turbulence in the outer disc.
@@
4.1.2 Stellar accretion in regions of low gas density
The second objection raised by Hartmann et al. (2006) is that stars in ionised HII regions still accrete. The typical density in such regions is far lower, and the sound speed and/or velocity dispersion much larger, such that the BHL accretion rate may be suppressed by several orders of magnitude. The example quoted by Hartmann et al. (2006) are some stars in Trumpler 37 which were reported by Sicilia-Aguilar et al. (2005) to be accreting substantially at a rate $\sim 10^{-8}\,M_{\odot}$ yr-1. This is despite a sound speed $c_{\mathrm{s,ext}}\sim 10$ km s-1 (which we now distinguish from the sound speed in the disc $c_{\mathrm{s,disc}}$) and an electron density $N_{\mathrm{e}}\sim 3{-}15$ cm-3 or $\rho_{\mathrm{gas}}\sim 2\times 10^{-23}$ g cm-3. The corresponding BHL accretion rate would therefore be $\dot{M}_{\mathrm{BHL}}\sim 10^{-13}\,M_{\odot}$ yr-1 for a solar mass star, far too small to drive observed accretion rates. This would appear to rule out in-falling gas as the origin of angular momentum transport in at least these discs. However, this conclusion requires further investigation for the following reasons.
Firstly, as discussed above, turbulence injected into the outer disc may only serve to feed internally mediated transport in the inner disc. Therefore turning off the supply from the reservoir of material accreted from the ISM should not immediately end accretion onto the star.
Secondly, the small BHL radius resulting from large $c_{\mathrm{s,ext}}$ means that, at least initially after the gas is ionised, accretion onto the disc will be facilitated by sweep-up and not BHL accretion (see Appendix G.3). If we re-calculate the sweep-up rate for a disc outer radius $250$ au (expected for a solar mass star), this increases the gas capture rate by two orders of magnitude to $\dot{M}_{\mathrm{su}}\sim 10^{-11}\,M_{\odot}$ yr-1. While this remains much smaller than the accretion rates reported by Sicilia-Aguilar et al. (2005) in Trumpler 37, if the HII region is young ($\lesssim 0.1-1$ Myr – Mellema et al., 2006; Raycheva et al., 2022), then there may still be significant density variations that can lead to a spread in BHL capture rates. Indeed, Sicilia-Aguilar et al. (2005) point out a population of embedded stars in Trumpler 37, which are evidence for at least some residual cold, dense gas.
Finally, as the mode of mass accretion changes, so does the rate of turbulent energy injection. This is both because the hot turbulent medium will typically have larger relative velocity $\Delta v_{\mathrm{gas}}\sim c_{\mathrm{s,ext}}$, and the material also injects thermal energy. In addition, if turbulent fluctuations are suppressed in the ionised medium, then accreted material contains little angular momentum and can fall directly into the inner disc regions. To estimate the resultant turbulent velocity within the disc, we can assume that $v_{\mathrm{in}}=\zeta c_{\mathrm{s,ext}}$, where $\zeta$ is a factor of order unity. Then:
$\dot{E}_{\mathrm{in}}=\frac{1}{2}\dot{M}_{\mathrm{su}}\zeta^{2}c_{\mathrm{s,ext}}^{2},$ (19)
where $\dot{M}_{\mathrm{su}}$ is the sweep-up rate. The resultant steady state Mach number is:
$\frac{v_{\mathrm{t,disc}}}{c_{\mathrm{s,disc}}}\approx 0.016\epsilon_{\mathrm{t}}^{1/3}\left(\frac{H_{\mathrm{disc}}}{10\,\mathrm{au}}\right)^{1/3}\left(\frac{M_{\mathrm{disc}}}{10^{-2}\,M_{\odot}}\right)^{-1/3}\left(\frac{\dot{M}_{\mathrm{su}}}{10^{-11}\,M_{\odot}\,\mathrm{yr}^{-1}}\right)^{1/3}\left(\frac{\zeta c_{\mathrm{s,ext}}}{10\,\mathrm{km}\,\mathrm{s}^{-1}}\right)^{2/3}\left(\frac{c_{\mathrm{s,disc}}}{0.2\,\mathrm{km}\,\mathrm{s}^{-1}}\right)^{-1}$ (20)
While for our chosen normalisation yields $\alpha_{\mathrm{SS}}\sim 10^{-4}$, some combination of a factor three increase in the in-fall velocity or a factor eight in the sweep-up accretion rate would be sufficient to drive more substantial $\alpha_{\mathrm{SS}}\sim 10^{-3}$. Turbulence injected by warping could also still play a role in this context (Deng et al., 2020). It is questionable whether either directly injected or warp-induced turbulence can sustain accretion into the inner disc for long in an ionised medium, but the moderate factors involved motivate further numerical experiment.
4.1.3 A case study: DM Tau
@@
We can also consider how the BHL accretion scenario fits with observations of individual discs. We consider the interesting scase of DM Tau, which is a star of mass $\sim 0.4{-}0.5\,M_{\odot}$ hosting a disc with a dust cavity and accreting at a rate $\sim 10^{-8}\,M_{\odot}$ yr-1 (Manara et al., 2014; Francis & van der Marel, 2020). It has a turbulent velocity $v_{\rm{t,disc}}\sim 0.25{-}0.33\,c_{\mathrm{s}}$, or viscous $\alpha_{\mathrm{SS}}\sim 0.1$ (Flaherty et al., 2020). For this disc, a recent and relatively rapid period of accretion would be needed to drive the observed turbulence. Estimating $M_{\mathrm{disc}}/m_{*}\approx 0.03$ (Andrews et al., 2013) this would imply recent accretion from the ISM at a rate $\dot{M}_{\mathrm{BHL}}\gtrsim 2\times 10^{-8}\,M_{\odot}$ yr-1 that ended in the last $\sim 10^{4}$ years. This is similar to the current stellar accretion rate (Manara et al., 2014), suggesting that the majority of the mass of the present day disc could have originated from this putative in-fall event. This scenario is plausible statistically within the context of our model. However, apart from the argument above, there is currently no particular evidence that DM Tau has undergone recent in-fall. This underlines the importance of a systematic search for observational signatures of late stage in-fall events.
4.1.4 To what degree does in-fall mediate angular momentum transport?
In this section, we have considered whether in-fall can drive the separate but related phenomena of turbulence and stellar accretion in protoplanetary discs. It is clear that there is sufficient energy in the in-falling material to drive turbulence compatible with what is observed in the surface layers of the outer disc in the few cases where such constraints exist. This is not the same as the statement that this in-fall does drive observed turbulence, to conclude as such will require careful case studies of observational examples. Similarly, understanding whether in-fall can regulate stellar accretion, either by angular momentum extraction or by turbulent energy injection, requires both more detailed theoretical and observational analysis. Probably internal processes must be responsible for angular momentum transport in at least some regions of the disc, but in-fall could play a role in regulating this process – for example, by supplying material from the outer disc. We conclude that, although it is far from clear that these fundamental problems in disc evolution are solved by in-fall, ruling out its importance is not presently justified.
4.2 Appearance of BHL-accreting discs
A pertinent question is what we expect the appearance of a disc undergoing BHL accretion to be. As previously discussed, when accreting material with with a high vorticity, the resultant disc outer radius can be of order the BHL radius (Krumholz et al., 2006; Kuffmeier et al., 2023). What is more uncertain is what should be the structure of a disc if the angular momentum of the accreted material is not aligned with the existing material. Depending on the relative angular momentum vector, this could result in ‘puffed up’ discs, or misalignment between the inner and outer regions (Kuffmeier et al., 2021). By contrast, at least the mm-dust component of discs is frequently observed to be extremely settled (Duchêne et al., 2003; Pinte et al., 2007, 2016; Villenave et al., 2020, 2022; Pizzati et al., 2023). However, (sub-)mm emission is also commonly more centrally concentrated than the gaseous disc by a factor $\sim 2.5$ (Ansdell et al., 2018; Facchini et al., 2019; Andrews, 2020), which may be due to radial drift or optical effects (Trapman et al., 2019; Rosotti et al., 2019). It is not clear whether turbulence driven in the outer disc should be reflected in the dust distribution, that may have drifted inwards and/or settled vertically, even in the presence of the warping induced by in-fall (Aly et al., 2024). In fact, we can consider one of the best known discs, HL Tau (ALMA Partnership et al., 2015), which has an extremely settled dust disc, evidenced by both dust height and polarisation that suggest low $\alpha_{\mathrm{SS}}\sim 10^{-5}$ (Pinte et al., 2016; Ueda et al., 2021). It is also one of the earliest known examples of a disc being fed by in-falling gas (Hayashi et al., 1993; Welch et al., 2000; Yen et al., 2019; Garufi et al., 2022) at a substantial rate ($\gtrsim 5\times 10^{-9}\,M_{\odot}$ yr-1 – Gupta et al., 2024). This example therefore makes a compelling argument that the mm dust component of the disc is not necessarily strongly influenced by in-fall. It also suggests that turbulence may be substantially lower in the mid-plane with respect to the surface layers. This represents an important topic for future study.
4.3 Observability of in-fall events
@@
where $\epsilon_{\mathrm{geo}}\leq 1$ is a geometric factor accounting for the non-uniform distribution of material and $\mu$ is the mean molecular mass. We will assume $\epsilon_{\mathrm{geo}}=1$ here, although based on the diverse ambient structures seen in observations, a wide range of $\epsilon_{\mathrm{geo}}$ may be plausible (e.g. Garufi et al., 2024; Zurlo et al., 2024).
@@
The typical column density implied by equation 22 is relatively low, such that we would expect that in-fall is often challenging to trace observationally. We may consider the example of CO isotopologues, which are often the brightest tracers used to probe the gas content of protoplanetary discs with ALMA. These isotopologues have a self-shielding column density of $\sim 10^{15}$ cm-2 (van Dishoeck & Black, 1988; Visser et al., 2009), and typical abundance relative to molecular hydrogen $\sim 10^{-5}$ (Trapman et al., 2023). As an order of magnitude estimate, we may therefore expect column densities $\lesssim 10^{20}$ cm-2 to have the CO photodissociated. From equation 22, we plot the $N_{\mathrm{ext}}$ distribution over time resulting from the accretion rates in our model in Figure 9. We see immediately in-falling CO would be photodissociated for the majority of discs in the age range $\sim 1{-}3$ Myr. Interestingly, in this age range, the fraction of discs above the threshold $N_{\mathrm{ext}}=10^{20}$ cm-2, approximately the CO self-shielding column, drops from approximately $30$ percent to close to 00 percent. Given that this is the range of ages in Taurus, this would be commensurate with the inferred $\sim 16$ percent of discs with evidence of interaction with the ambient medium from scattered light observations (Garufi et al., 2024). However, the physical column density down to which these observations are sensitive is not clear. While these estimates are approximate, they highlight the importance of deep surveys of the environments surrounding large numbers of discs. Such data could be applied to make statistical inferences about the prevalence of late stage in-fall, correlating it with disc properties and testing the theoretical predictions we have made in this work.